# Large-Scale Continual Scheduling and Execution for Dynamic Distributed Satellite Constellation Observation Allocation

### v1

@@ Abstract.
The size and capabilities of Earth-observing satellite constellations are rapidly increasing. Leveraging distributed onboard control, we can enable novel time-sensitive measurements and responses. However, deploying autonomy to satellites requires efficient computation and communication. This work tackles the challenge of efficiently scheduling observations for hundreds of satellites in a dynamic, large-scale problem with millions of variables. We present the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP), a new formulation of Dynamic Distributed Constraint Optimization Problems (DDCOP) that models integrated scheduling and execution. DCOSP has a novel optimality condition for which we construct an omniscient offline algorithm for its computation. We also present the Dynamic Incremental Neighborhood Stochastic Search algorithm (D-NSS), an incomplete online decomposition-based DDCOP algorithm that repairs and solves sub-problems when problem dynamics occur. We show through simulation that D-NSS converges to near-optimal solutions and outperforms DDCOP baselines in terms of solution quality, computation time, and message volume. As part of the NASA FAME mission, DCOSP and D-NSS will be the foundation of the largest in-space demonstration of distributed multi-agent AI to date.
@@ 1. Introduction
There has been a proliferation of Earth-observing spacecraft in recent years, including advancements in their capabilities to act as autonomous agents. Reduced launch costs have led to constellations composed of hundreds or thousands of spacecraft that can monitor Earth phenomena NewSpace (2023). Large observation systems result in shorter revisit times to the target observation locations. Reduced revisit times are crucial for rapid responses to dynamic events such as natural disasters. New spacecraft also possess edge hardware capable of performing more intensive computation onboard, including neural network execution and even planning Chien et al. (2005); Chien et al. (2024); Zilberstein et al. (2024); Chien et al. (2025b); Rabideau et al. (2025). The advancement of inter-satellite links (ISL) has enabled persistent communications between spacecraft and stations on the ground.
@@
These capabilities support observation campaigns that require time-sensitive and coordinated measurements. An example is global monitoring of all volcanic activity or flooding events. However, without consistent observation and the ability to react to dynamic events, key measurements of these processes may be missed. For example, wildfire monitoring requires 30-minute updates to be useful to ground responders Kopardekar and Grindle (2021). Centralized scheduling on the ground suffers from latency and may not be able to meet these timing constraints. Therefore, to successfully tackle these campaigns, we require large-scale observation systems with reduced revisit times and distributed autonomy so that agents can operate online without ground control.
@@
This work focuses on distributed online scheduling that can efficiently coordinate the actions of hundreds or thousands of spacecraft continually as problem dynamics change. Coordinating the actions of a large-scale constellation requires reasoning about agents with varying capabilities, constraints, and visibility of Earth targets while managing limited computational resources. Satellites share CPU and RAM with other critical flight software. Large volumes of communication in space are also unreliable and may even carry a financial cost Lin et al. (2024). These limitations make solving a static problem challenging. However, any operational solution must solve dynamic problems that change over time, which further increases the complexity. In these scenarios, the satellite constellation is fixed, yet the observation requests change. These changing observation requests alter the constraints of the problem. We desire continual scheduling solutions that are lightweight yet can effectively handle these problem dynamics.
@@
Operational satellite observation scheduling uses centralized paradigms Shah et al. (2019), and many research efforts have concentrated on centralized solutions Boerkoel et al. (2021); Nag et al. (2018). Centralized approaches may be insufficient for dynamic situations that require real-time response due to latencies to the ground. Centralized approaches are also vulnerable to single-point failures that would result in a non-operational constellation.
@@
DCOSP problem instances are much larger than typical DDCOP problems examined in previous literature. DCOSP instances consist of millions of variables that change over time. Agents must react to problem dynamics while expending limited time, memory, and communication. We extend the Neighborhood Stochastic Search algorithm (NSS) to a dynamic variation referred to as Dynamic Incremental Neighborhood Stochastic Search (D-NSS). NSS decomposes the global problem into smaller sub-problems, producing global solutions more efficiently. D-NSS uses the same foundation and leverages repairing previous solutions to efficiently handle problem dynamics.
@@
    a novel formulation of the real-world application of dynamic satellite scheduling as a DDCOP with a unique optimality condition that models task execution,
@@
We evaluate the effectiveness of the approaches on large-scale real-world scenarios as well as analyze the challenges of deploying existing DDCOP solutions to this problem. DCOSP and D-NSS will be leveraged in the largest in-space demonstration of multi-agent AI to date, beginning in 2026. The NASA FAME demonstration involves over 60 participating spacecraft that will dynamically coordinate to observe Earth phenomena Chien et al. (2025a).
@@ 2.1. Satellite Observation Scheduling
Satellite observation scheduling is typically framed as an optimization problem that involves geometric reasoning, downlink scheduling, and constraint-based task allocation. The majority of research efforts and operational work have focused on centralized solutions to satellite observation scheduling Globus et al. (2004); Augenstein et al. (2016); Nag et al. (2018); He et al. (2018); Shah et al. (2019); Wang et al. (2020); Squillaci et al. (2021); Boerkoel et al. (2021); Eddy and Kochenderfer (2021); Squillaci et al. (2023); Chatterjee and Tharmarasa (2024). There is limited work on decentralized scheduling approaches, and these mainly focus on static problems. Examples include auction-based methods Picard (2022); Phillips and Parra (2021) and heuristic search-based methods Parjan and Chien (2023); Bonnet and Tessier (2007, 2008); Zilberstein et al. (2025b). The work of Zilberstein, Rao, Salis, and Chien proposed the Multi-Satellite Constellation Observation Scheduling Problem (COSP), a DCOP formulation of observation allocation Zilberstein et al. (2025b). We extend this work by formulating the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP), which is a novel DDCOP formulation of the problem.
@@ 2.2. Dynamic Distributed Constraint Optimization
Distributed constraint optimization problems (DCOP) have modeled applications including mobile sensor teams Pertzovsky et al. (2024), smart grids Fioretto et al. (2017), and satellite scheduling Zilberstein et al. (2025b). Solutions to distributed constraint optimization problems tend to be intensive in computation and communication, making deployment to agents with limited computation challenging. Optimal solutions have exponential complexities Hirayama and Yokoo (1997); Modi et al. (2005); Gershman et al. (2009); Petcu and Faltings (2005); Mailler and Lesser (2004). Incomplete DCOP algorithms are more efficient, yet typically rely on agents communicating with all neighboring agents in the constraint graph, resulting in large complexities when constraint graphs are fully connected Maheswaran et al. (2004); Zhang et al. (2005); Stranders et al. (2009); Nguyen et al. (2019). The Neighborhood Stochastic Search (NSS) algorithm, which iteratively improves sub-problem solutions, has been shown to solve large-scale distributed satellite observation scheduling with limited computation and communication Zilberstein et al. (2025b). We extend the NSS algorithm to the dynamic DCOP setting (DDCOP) to perform scalable and effective dynamic observation scheduling.
Dynamic distributed constraint optimization problems (DDCOP) Lass et al. (2008) extend DCOPs to capture problem changes. A standard DDCOP is composed of a sequence of $T$ static DCOPs where an optimal solution is obtained by solving each of the $T$ DCOPs optimally. DDCOP solutions are inherently online algorithms as a system reacts to changes. Most work has adapted common DCOP algorithms to dynamic variations that inherit computational complexities Mailler (2005); Faltings (2005); Khanna et al. (2009); Billiau et al. (2010); Zivan et al. (2015); Yeoh et al. (2015); Ridgway and Mailler (2015).
@@ 3.1. Multi-Satellite Constellation Observation Scheduling
COSP has been shown to be a challenging problem for DCOP methods. Typical COSP instances have hundreds of agents and thousands of requests, leading to millions of decision variables. In addition, the constraint graph of COSP has high degrees on the order of $\Omega(|A|\cdot|R|)$. The constraint graph is also assumed to be only locally known to an agent. An agent $a$ is oblivious to all tasks $s\notin S_{a}$ and variables $x\notin\mathcal{X}_{a}$. This is not consistent with standard DCOPs in which agents know all neighboring variables/agents in the constraint graph Fioretto et al. (2018). In COSP, it is assumed that agents know the existence of all other agents but have no knowledge of the variables of other agents. These factors make existing DCOP approaches that rely on agents communicating with neighboring agents in the constraint graph both computationally challenging due to the high degrees in the graph and inapplicable since we cannot assume agents know which agents they share constraints with.
@@ 4.1. Obtaining an Optimal Solution to DCOSP
Solving $\delta^{\prime}$ optimally results in an optimal solution to the DCOSP $\delta$. By definition, $\delta^{\prime}$ is a static COSP instance. Therefore, we can employ any complete algorithm to obtain an optimal solution to $\delta$ by solving $\delta^{\prime}$. Although the omniscient solver is not deployable, it serves as a key performance upper bound for benchmarking practical, online algorithms such as D-NSS. However, even solving a single DCOP is NP-Hard Modi et al. (2005). Therefore, for large DCOSP instances, we use an incomplete solver to obtain a near-optimal solution since any exponential time search is infeasible.
@@ 4.2. Dynamic Incremental Neighborhood Stochastic Search
In this section, we present the Dynamic Incremental Neighborhood Stochastic Search algorithm (D-NSS) shown in Algorithm 1. Due to the scale of DCOSP and the computational constraints of satellites, we require DDCOP algorithms that are both efficient and can reason about the dynamic nature of the problem. D-NSS extends NSS to address the drawbacks of general DDCOP algorithms. NSS is an iterative algorithm where at each iteration agents stochastically update their variable assignments based on the assignments of agents they communicate with. This iterative procedure is shared with other algorithms such as DSA. However, NSS relies on a decomposition heuristic $\Upsilon:A\times S\rightarrow\{0,1\}$ to generate a subproblem $\mathcal{N}$ for neighborhoods of agents to solve. This subproblem is a smaller DCOP consisting of requests $R_{\mathcal{N}}$ and agents $A_{\mathcal{N}}$. D-NSS continually computes subproblems, repairs local solutions between problem instances, and reasons about prior scheduling and execution in the search and repair phases.
@@
We leverage the Geometric Neighborhood Decomposition heuristic (GND) to create these subproblems every time dynamics occur Zilberstein et al. (2025b). GND efficiently allocates requests to neighborhoods of agents based on the orbital geometry of the constellation and has been shown to effectively partition COSP instances. For completeness, we provide an outline of GND. GND was first introduced with the NSS algorithm and we refer the reader to prior work for a full presentation of the heuristic Zilberstein et al. (2025b).
GND leverages the orbital geometry of the satellite constellation to hierarchically partition requests to neighborhoods of agents. Let $K$ be the set of orbital planes that define a satellite constellation [^1] An orbital plane is a fixed orbit around Earth that many satellites may follow at various spacing.. For every request, an agent computes the supply from all orbital planes and adds these to estimate the total number of agents with overflights for the request. The supply can also be thought of as an estimate of degrees in the constraint graph. Iterating through requests in ascending order of supply, a request gets assigned to the $n$ neighborhoods with the highest ratio of supply to temporal conflicts. Temporal conflicts are counts of other requests already allocated to a neighborhood that overlap in time with a given request. Finally, within a neighborhood, requests are further subdivided to agents based on biases towards specific tiles on Earth. GND with $n$ degrees of incompleteness is denoted GND($n$).
@@
D-NSS restarts the search phase when changes are initiated in the problem. A key procedure is the repair function that repairs previously computed solutions to leverage assignments of unchanging variables shown in Algorithm 2. Repairing consists of removing all tasks that are no longer in the current problem instance and greedily inserting new tasks in random order. Note that we can also remove from an agent’s schedule all requests that have been previously executed by agents in the same neighborhood. One benefit of the D-NSS algorithm is that it can leverage different repair procedures. We use the random repair function for two main reasons. Random initialization is the standard for DSA and NSS variants in static domains Zhang et al. (2005) as it promotes diverse solutions and D-NSS is designed to be as lightweight as possible. Computation-intensive repair procedures counteract the efficiency of the algorithm. We show later on that performing random repair improves the quality and efficiency of D-NSS on DCOSP instances.
@@ 4.3. Baseline Algorithms
We evaluate several baseline algorithms in addition to D-NSS. The naive alternative to D-NSS is to recompute a solution from scratch every time problem dynamics occur. We refer to this procedure as 0-NSS. 0-NSS has the same theoretical complexities as D-NSS. We also evaluate dynamic variations of the Distributed Stochastic Search Algorithm (DSA) Zhang et al. (2005); Parjan and Chien (2023). D-DSA uses Algorithm 2 to repair DSA solutions. Likewise, 0-DSA runs DSA from scratch every time the problem changes. These DSA variants have a complexity of $O(|A|{\cdot}|R|)$ per iteration. Dynamic DSA is one of the most lightweight DDCOP solvers and can be deployed without violating the constraint graph assumptions of DCOSP. Other DDCOP algorithms such as a dynamic variation of MGM Maheswaran et al. (2004), would also incur a computation and communication complexity of $O(|A|{\cdot}|R|)$ per iteration. Due to the scale of the problem instances and the constraints of satellite computation and communication, we are unable to evaluate any exponential-time algorithms.
We include two baseline non-communication-reliant algorithms: random and greedy. Both of these algorithms construct a schedule for a satellite without reasoning about other agents. These algorithms transition schedules after the dynamics occur by removing redundant tasks. The random solver randomly inserts tasks into an agent’s schedule while the greedy algorithm orders tasks by their start time and iteratively inserts them into a schedule in a single pass. Although simple, a greedy solver is comparable to deployed planners on spacecraft Gaines et al. (2022).
Finally, we use the construction outlined previously to obtain optimal and near-optimal solutions. For small problems, we obtain an optimal solution by constructing a static COSP instance and solving it with a centralized branch and bound. For large problems, computing an optimal solution is not computationally feasible, so we lower bound the optimal solution using Squeaky Wheel Optimization Joslin and Clements (1999), an incomplete centralized solver.

### v2

@@ Abstract.
The size and capabilities of Earth-observing satellite constellations are rapidly increasing. Leveraging distributed onboard control, we can enable novel time-sensitive measurements and responses. However, deploying autonomy to large multiagent satellite systems necessitates algorithms with efficient computation and communication. We tackle this challenge and propose new, online algorithms for large-scale dynamic distributed constraint optimization problems (DDCOP). We present the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP), a new formulation of DDCOPs that models integrated scheduling and execution. We construct an omniscient offline algorithm to compute the novel optimality condition of DCOSP and present the Dynamic Incremental Neighborhood Stochastic Search (D-NSS) algorithm, an incomplete online decomposition-based DDCOP approach. We show through simulation that D-NSS converges to near-optimal solutions and outperforms DDCOP baselines in terms of solution quality, computation time, and message volume. Our work forms the foundation of the largest in-space demonstration of distributed multiagent AI to date: the NASA FAME mission.
@@ 1. Introduction
There has been a proliferation of Earth-observing spacecraft in recent years, including advancements in their capabilities to act as autonomous agents. Reduced launch costs have led to constellations composed of hundreds or thousands of spacecraft that can monitor Earth phenomena newspace. Large observation systems result in shorter revisit times to the target observation locations. Reduced revisit times are crucial for rapid responses to dynamic events such as natural disasters. New spacecraft also possess edge hardware capable of performing more intensive computation onboard, including neural network execution and even planning eo1; nos2024; zilberstein-isairas-2024; dt-spaceops-2025; rabideau2025planning. The advancement of inter-satellite links (ISL) has enabled persistent communications between spacecraft and stations on the ground.
@@
These capabilities support observation campaigns that require time-sensitive and coordinated measurements. An example is global monitoring of all volcanic activity or flooding events. However, without consistent observation and the ability to react to dynamic events, key measurements of these processes may be missed. For example, wildfire monitoring requires 30-minute updates to be useful to ground responders wildfireworkshop. Centralized scheduling on the ground suffers from latency and may not be able to meet these timing constraints. Therefore, to successfully tackle these campaigns, we require large-scale observation systems with reduced revisit times and distributed autonomy so that agents can operate online without ground control.
@@
This work focuses on distributed online scheduling that can efficiently coordinate the actions of hundreds or thousands of spacecraft continually as problem dynamics change. Coordinating the actions of a large-scale constellation requires reasoning about agents with varying capabilities, constraints, and visibility of Earth targets while managing limited computational resources. Satellites share CPU and RAM with other critical flight software. Large volumes of communication in space are also unreliable and may even carry a financial cost leo-comms. These limitations make solving a static problem challenging. However, any operational solution must solve dynamic problems that change over time, which further increases the complexity. In these scenarios, the satellite constellation is fixed, yet the observation requests change. These changing observation requests alter the constraints of the problem. We desire continual scheduling solutions that are lightweight yet can effectively handle these problem dynamics.
@@
Operational satellite observation scheduling uses centralized paradigms shah2019scheduling, and many research efforts have concentrated on centralized solutions boerkoel2021efficientsatellite; nag2018scheduling. Centralized approaches may be insufficient for dynamic situations that require real-time response due to latencies to the ground. Centralized approaches are also vulnerable to single-point failures that would result in a non-operational constellation.
@@
DCOSP problem instances are much larger than typical DDCOP problems examined in previous literature. DCOSP instances consist of millions of variables that change over time. Agents must react to problem dynamics while expending limited time, memory, and communication. We extend the Neighborhood Stochastic Search (NSS) algorithm to a dynamic variation referred to as Dynamic Incremental Neighborhood Stochastic Search (D-NSS). NSS decomposes the global problem into smaller sub-problems, producing global solutions more efficiently. D-NSS uses the same foundation and leverages repairing previous solutions to efficiently handle problem dynamics.
@@
    formulating the real-world application of dynamic satellite scheduling as a DDCOP with a unique optimality condition that models task execution,
@@
We evaluate the effectiveness of the approaches on large-scale real-world scenarios as well as analyze the challenges of deploying existing DDCOP solutions to this problem. DCOSP and D-NSS will be leveraged in the largest in-space demonstration of multi-agent AI to date, beginning in 2026. The NASA FAME demonstration involves over 60 participating spacecraft that will dynamically coordinate to observe Earth phenomena fame-spaceops-2025.
@@ 2.1. Satellite Observation Scheduling
Satellite observation scheduling is typically framed as an optimization problem that involves geometric reasoning, downlink scheduling, and constraint-based task allocation. The majority of research efforts and operational work have focused on centralized solutions to satellite observation scheduling globus; augenstein2016optimal; nag2018scheduling; he2018improved; shah2019scheduling; aeossp; squillaci2021managing; boerkoel2021efficientsatellite; eddy2021maximum; squillaci2023scheduling; sat-task-plan-large-areas. There is limited work on decentralized scheduling approaches, and these mainly focus on static problems. Examples include auction-based methods picardauction; phillipsauction and heuristic search-based methods shreya; bonnet2007collaboration; bonnet2008coordination; zilberstein2025decentralized. The work of Zilberstein, Rao, Salis, and Chien proposed the Multi-Satellite Constellation Observation Scheduling Problem (COSP), a DCOP formulation of observation allocation zilberstein2025decentralized. We extend this work by formulating the Dynamic Multi-Satellite Constellation Observation Scheduling Problem (DCOSP), which is a novel DDCOP formulation of the problem.
@@ 2.2. Dynamic Distributed Constraint Optimization
Distributed constraint optimization problems (DCOP) have modeled applications including mobile sensor teams ca-dcop-mst, smart grids smartgrid, and satellite scheduling zilberstein2025decentralized. Solutions to distributed constraint optimization problems tend to be intensive in computation and communication, making deployment to agents with limited computation challenging. Optimal solutions have exponential complexities syncbb; adopt; gershman2009asynchronous; dpop; mailler2004solving. Incomplete DCOP algorithms are more efficient, yet typically rely on agents communicating with all neighboring agents in the constraint graph, resulting in large complexities when constraint graphs are fully connected mgm; dsa; maxsum; dgibs. The Neighborhood Stochastic Search (NSS) algorithm, which iteratively improves sub-problem solutions, has been shown to solve large-scale distributed satellite observation scheduling with limited computation and communication zilberstein2025decentralized. We extend the NSS algorithm to the dynamic DCOP setting (DDCOP) to perform scalable and effective dynamic observation scheduling.
Dynamic distributed constraint optimization problems (DDCOP) lass2008dynamic extend DCOPs to capture problem changes. A standard DDCOP is composed of a sequence of $T$ static DCOPs where an optimal solution is obtained by solving each of the $T$ DCOPs optimally. DDCOP solutions are inherently online algorithms as a system reacts to changes. Most work has adapted common DCOP algorithms to dynamic variations that inherit computational complexities mailler2005comparing; faltings2005superstabilizing; khanna2009efficient; billiau2012sbdo; zivan2015distributed; yeoh2015incremental; dynamicdsa.
@@ 3.1. Multi-Satellite Constellation Observation Scheduling
COSP has been shown to be a challenging problem for DCOP methods. Typical COSP instances have hundreds of agents and thousands of requests, leading to millions of decision variables. In addition, the constraint graph of COSP has high degrees on the order of $\Omega(|A|\cdot|R|)$. The constraint graph is also assumed to be only locally known to an agent. An agent $a$ is oblivious to all tasks $s\notin S_{a}$ and variables $x\notin\mathcal{X}_{a}$. This is not consistent with standard DCOPs in which agents know all neighboring variables/agents in the constraint graph fioretto2018distributed. In COSP, it is assumed that agents know the existence of all other agents but have no knowledge of the variables of other agents. These factors make existing DCOP approaches that rely on agents communicating with neighboring agents in the constraint graph both computationally challenging due to the high degrees in the graph and inapplicable since we cannot assume agents know which agents they share constraints with.
@@ 4.1. Obtaining an Optimal Solution to DCOSP
Solving $\delta^{\prime}$ optimally results in an optimal solution to the DCOSP $\delta$. By definition, $\delta^{\prime}$ is a static COSP instance. Therefore, we can employ any complete algorithm to obtain an optimal solution to $\delta$ by solving $\delta^{\prime}$. Although the omniscient solver is not deployable, it serves as a key performance upper bound for benchmarking practical, online algorithms such as D-NSS. However, even solving a single DCOP is NP-Hard adopt. Therefore, for large DCOSP instances, we use an incomplete solver to obtain a near-optimal solution since any exponential time search is infeasible.
@@ 4.2. Dynamic Incremental Neighborhood Stochastic Search
In this section, we present the Dynamic Incremental Neighborhood Stochastic Search (D-NSS) algorithm shown in Algorithm 1. Due to the scale of DCOSP and the computational constraints of satellites, we require DDCOP algorithms that are both efficient and can reason about the dynamic nature of the problem. D-NSS extends NSS to address the drawbacks of general DDCOP algorithms. NSS is an iterative algorithm where at each iteration agents stochastically update their variable assignments based on the assignments of agents they communicate with. This iterative procedure is shared with other algorithms such as DSA. However, NSS relies on a decomposition heuristic $\Upsilon:A\times S\rightarrow\{0,1\}$ to generate a subproblem $\mathcal{N}$ for neighborhoods of agents to solve. This subproblem is a smaller DCOP consisting of requests $R_{\mathcal{N}}$ and agents $A_{\mathcal{N}}$. D-NSS continually computes subproblems, repairs local solutions between problem instances, and reasons about prior scheduling and execution in the search and repair phases.
@@
We leverage the Geometric Neighborhood Decomposition (GND) heuristic to create these subproblems every time dynamics occur zilberstein2025decentralized. GND efficiently allocates requests to neighborhoods of agents based on the geometry of the constellation and has been shown to effectively partition COSP instances. For completeness, we provide an outline of GND. GND was first introduced with the NSS algorithm and we refer the reader to prior work for a full presentation of the heuristic zilberstein2025decentralized.
GND leverages the orbital geometry of the satellite constellation to hierarchically partition requests to neighborhoods of agents. Let $K$ be the set of orbital planes that define a satellite constellation. For every request, an agent computes the supply from all orbital planes and adds these to estimate the total number of agents with overflights for the request. The supply can also be thought of as an estimate of degrees in the constraint graph. Iterating through requests in ascending order of supply, a request gets assigned to the $n$ neighborhoods with the highest ratio of supply to temporal conflicts. Temporal conflicts are counts of other requests already allocated to a neighborhood that overlap in time with a given request. Finally, within a neighborhood, requests are further subdivided to agents based on biases towards specific tiles on Earth. GND with $n$ degrees of incompleteness is denoted GND($n$).
@@
D-NSS restarts the search phase when changes are initiated in the problem. A key procedure is the repair function that repairs previously computed solutions to leverage assignments of unchanging variables shown in Algorithm 2. Repairing consists of removing all tasks that are no longer in the current problem instance and greedily inserting new tasks in random order. Note that we can also remove from an agent’s schedule all requests that have been previously executed by agents in the same neighborhood. One benefit of the D-NSS algorithm is that it can leverage different repair procedures. We use the random repair function for two main reasons. Random initialization is the standard for DSA and NSS variants in static domains dsa as it promotes diverse solutions and D-NSS is designed to be as lightweight as possible. Computation-intensive repair procedures counteract the efficiency of the algorithm. We show later on that performing random repair improves the quality and efficiency of D-NSS on DCOSP instances.
@@ 4.3. Baseline Algorithms
We evaluate several baseline algorithms in addition to D-NSS. The naive alternative to D-NSS is to recompute a solution from scratch every time problem dynamics occur. We refer to this procedure as 0-NSS. 0-NSS has the same theoretical complexities as D-NSS. We also evaluate dynamic variations of the Distributed Stochastic Search (DSA) algorithm dsa; shreya. D-DSA uses Algorithm 2 to repair DSA solutions. Likewise, 0-DSA runs DSA from scratch every time the problem changes. These DSA variants have a complexity of $O(|A|{\cdot}|R|)$ per iteration. Dynamic DSA is one of the most lightweight DDCOP solvers and can be deployed without violating the constraint graph assumptions of DCOSP. Other DDCOP algorithms such as a dynamic variation of MGM mgm, would also incur a computation and communication complexity of $O(|A|{\cdot}|R|)$ per iteration. Due to the scale of the problem instances and the constraints of satellite computation and communication, we are unable to evaluate any exponential-time algorithms.
We include two baseline non-communication-reliant algorithms: random and greedy. Both of these algorithms construct a schedule for a satellite without reasoning about other agents. These algorithms transition schedules after the dynamics occur by removing redundant tasks. The random solver randomly inserts tasks into an agent’s schedule while the greedy algorithm orders tasks by their start time and iteratively inserts them into a schedule in a single pass. Although simple, a greedy solver is comparable to deployed planners on spacecraft m2020-planner-astra2022.
Finally, we use the construction outlined previously to obtain optimal and near-optimal solutions. For small problems, we obtain an optimal solution by constructing a static COSP instance and solving it with a centralized branch and bound. For large problems, computing an optimal solution is not computationally feasible, so we lower bound the optimal solution using Squeaky Wheel Optimization (SWO) swo, an incomplete centralized solver.